OpenStudy (anonymous):

Elena is thinking about putting $200 in a savings account that earns 4% interest compounded semiannually. She wants to keep that money in the account for 4 years. Which of the formulas below can help her calculate how much money she will have at the end of the 4 years? A. $200(1 + 0.02)8 B. $200(1 + 0.04)4 C. $200(1 + 0.02)4 D. $200(1 + 0.08)2

OpenStudy (anonymous):

@e.mccormick help please

OpenStudy (e.mccormick):

Have you seen the formula for compound interest?

OpenStudy (anonymous):

no sadly @e.mccormick

OpenStudy (jdoe0001):

\(\bf A=P\left(1+\frac{r}{n}\right)^{nt} \\ \quad \\ A=\textit{current amount}\\ P=\textit{original amount deposited}\to \$200\\ r=rate\to 4\%\to \frac{4}{100}\to 0.04\\ n=\textit{times it compounds per year, semi-annually}\to 2\\ t=years\to 4\)

OpenStudy (anonymous):

ohhhhh ok now what @jdoe0001 and @e.mccormick

OpenStudy (e.mccormick):

Oh, and here I was writing that thing up again. Hehe.

OpenStudy (e.mccormick):

You put in the numbers you were given for the letters and then solve.

OpenStudy (jdoe0001):

well.. plug and chug

OpenStudy (e.mccormick):

Well, solve partially. You just need to get it to the form used in your answers.

OpenStudy (anonymous):

ughhhhhhh i hate mathe i keep geeti crazy numbers @e.mccormick

OpenStudy (e.mccormick):

When I look at your answers I see two things that are sort of unique: the interest part and the rt part. Each one is only duplicated once. So, if you find either of those you will either eliminate half the answers or find the correct answer.

OpenStudy (anonymous):

ok thanks @e.mccormick